OpenStudy (anonymous):

How can consensus theorem be applied to prove A'B+B'C+C'A=AB'+BC'+CA'? Guys, please help me with this prob ASAP... Thnx in advance....

OpenStudy (anonymous):

A'B+B'C+C'A=A'B(C+C')+(A+A')B'C+A(B+B')C'= A'BC+A'BC'+AB'C+A'B'C+ABC'+AB'C'=AB'(C+C')+(A+A')BC'+A'(B+B')C= AB'+BC'+A'C

OpenStudy (anonymous):

Keep in mind that A+A'=1; B+B'=1;C+C'=1

OpenStudy (anonymous):

Thnx Carlos...